# Simplifying, Stabilizing & Scaling Continuous-Time Consistency Models

### v1

@@ B.2 Relationship with other parameterization
Flow matching (Lipman et al., 2022; Liu et al., 2022; Albergo et al., 2023) defines a stochastic path between two samples ${\bm{x}}_{0}$ from data distribution and ${\bm{z}}$ from a tractable distribution which is usually some Gaussian distribution. For a general path ${\bm{x}}_{t}=\alpha_{t}{\bm{x}}_{0}+\sigma_{t}{\bm{z}}$ with $\alpha_{0}=1,\alpha_{T}=0,\sigma_{0}=0,\sigma_{T}=1$, the conditional probability path is
@@ Appendix D Adaptive Weighting For Diffusion Models, Consistency Models and Variational Score Distillation
$\mathcal{L}_{\text{CM}}(\theta)=\nabla_{\theta}\mathbb{E}_{t,{\bm{x}}_{0},{\bm{z}}}\left[-w(t)\sin(t){\bm{f}}_{\theta}^{\top}\frac{\mathrm{d}{\bm{f}}_{\theta^{-}}}{\mathrm{d}t}\right],$
@@ E.1 Parameterization and Training Objective
     ${\bm{g}}\leftarrow-\cos^{2}(t)(\sigma_{d}{\bm{F}}_{\theta^{-}}-\sigma_{d}{\bm{F}}_{\text{pretrain}})-r\cdot\cos(t)\sin(t)({\bm{x}}_{t}+\sigma_{d}\frac{\mathrm{d}{\bm{F}}_{\theta^{-}}}{\mathrm{d}t})$triangleright JVP rearrangement
@@ G.2 Continuous-time Consistency Models
VSD experiments. We do not use EMA for ${\bm{F}}_{\phi}$ in VSD, instead we always use the original model for ${\bm{F}}_{\phi}$ for stabilizing the training. The learning rate of ${\bm{F}}_{\phi}$ is the same as the learning rate of CMs. More details and results are provided in Tables 4, 5 and 6.

### v2

@@ B.2 Relationship with other parameterization
Flow matching (Lipman et al., 2022; Liu et al., 2022; Albergo et al., 2023; Kornilov et al., 2024) defines a stochastic path between two samples ${\bm{x}}_{0}$ from data distribution and ${\bm{z}}$ from a tractable distribution which is usually some Gaussian distribution. For a general path ${\bm{x}}_{t}=\alpha_{t}{\bm{x}}_{0}+\sigma_{t}{\bm{z}}$ with $\alpha_{0}=1,\alpha_{T}=0,\sigma_{0}=0,\sigma_{T}=1$, the conditional probability path is
@@ Appendix D Adaptive Weighting For Diffusion Models, Consistency Models and Variational Score Distillation
$\nabla_{\theta}\mathcal{L}_{\text{CM}}(\theta)=\nabla_{\theta}\mathbb{E}_{t,{\bm{x}}_{0},{\bm{z}}}\left[-w(t)\sin(t){\bm{f}}_{\theta}^{\top}\frac{\mathrm{d}{\bm{f}}_{\theta^{-}}}{\mathrm{d}t}\right],$
@@ E.1 Parameterization and Training Objective
     ${\bm{g}}\leftarrow\cos(t)\Delta_{\theta^{-}}({\bm{x}}_{t},t,t^{\prime},r)$triangleright JVP rearrangement
@@ E.2 Experiments of Discrete-time sCM
E.3 Comparison with ECT
We compare the 1-step sampling FID scores at different training iterations between ECT (Geng et al., 2024) and sCT on CIFAR-10. As shown in Table 3, our proposed sCT significantly outperforms ECT during the training, demonstrating the effectiveness of the compute efficiency and faster convergence of sCT.
For fair comparison, we use the same network architecture with ECT on CIFAR-10, which is the DDPM++ network proposed by Ho et al. (2020) and does not have AdaGN layer, and use the same dropout rate of 0.20 as ECT, and use the same batch size (128) as ECT (which is different from our default setting of 512 in our reported results in Table 1). We choose $P_{\text{mean}}=-1.0$ and $P_{\text{std}}=1.8$ for sCT, and use TrigFlow parameterization (with $c_{\text{noise}}=t$). All the other hyperparameters are the same as the experiments in Table 1.
Table 3: Sample quality measured by FID score ($\downarrow$) of ECT (Geng et al., 2024) and sCT at different training iterations on CIFAR-10.
Training Iterations 100k 200k 400k
ECT 4.54 3.86 3.60
sCT (ours) 3.97 3.51 3.09
@@ G.2 Continuous-time Consistency Models
VSD experiments. We do not use EMA for ${\bm{F}}_{\phi}$ in VSD, instead we always use the original model for ${\bm{F}}_{\phi}$ for stabilizing the training. The learning rate of ${\bm{F}}_{\phi}$ is the same as the learning rate of CMs. More details and results are provided in Tables 5, 6 and 7.